# SpeechRole: A Large-Scale Dataset and Benchmark for Evaluating Speech Role-Playing Agents

### v3

@@ 1 Introduction
Recently, Role-Playing Agents (RPAs) have emerged as a promising paradigm for achieving personalized interactions and emotional resonance [8]. Simulated characters with distinct traits make interactions more engaging in applications such as digital assistants [43], educational tools [44], and interactive storytelling [36]. However, current research predominantly focuses on the text modality [23, 32, 36, 31], overlooking the crucial role of speech in authentic interactions.
Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [5]. Although preliminary studies have explored role-playing incorporating visual or speech modalities [11, 46], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs). SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions. In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
@@ 2 Related Work
Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality. ChatHaruhi [23] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [36], Character-LLM [32], and RoleFact [31] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations. Beyond text, MMRole [11] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [46] demonstrated the value of speech for real-time, voice-aware character interaction.
To assess role-playing capabilities, benchmarks such as RoleEval [33], RoleInteract [6], SocialBench [7], and CharacterEval [34] evaluated behavioral consistency and persona fidelity, while INCHARACTER [39] and ECHO [24] introduced human-aligned, psychometric evaluation paradigms. These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity. However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
@@
In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [21]. Modern TTS models like F5TTS [10], FireRedTTS [19], E2-TTS [16], Spark-TTS [38], CosyVoice2 [14], MaskGCT [40], and XTTS [4] have significantly improved output quality, expressiveness, and controllability. However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
End-to-end systems aim to unify speech perception and generation within a single model. Early efforts like SpeechGPT [45] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [29] utilized vocoder-based representations for continuous speech modeling. Recent advances, including Moshi [13], LLaMA-Omni [17], and Qwen2.5-Omni [41], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
@@ 3.1 Text Data Construction
Using the extracted metadata as a guide, we employ gpt-4.1-2025-04-14 [25] to generate conversations. For each of the 98 roles, we produce approximately 800 single-turn and 800 multi-turn dialogues, creating a substantial initial pool of text data.
@@
The generated data undergoes a multi-stage filtering process. First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [28] for English and text2vec-bge-large-chinese [42] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese. To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
@@ 3.2 Role Voice Collection
To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works. All audio is anonymized and fragmented into brief utterances to comply with fair use. Raw streams are extracted using ffmpeg [18] and converted to mono 16 kHz WAV format.
@@
We use the open-source Emilia framework [20] to process raw audio, including source separation, speaker diarization, and voice activity detection. This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [26] and scored with DNSMOS P.835 OVRL [27].
@@
The diarization process outputs speaker indices without explicit identity information. To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [12]) to analyze transcribed text. The models’ integrated results are then manually verified.
@@
To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity. We use the CAM++ model [35] to extract a speaker embedding for each clip. The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
@@ 4.3 Evaluation Procedure
There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [37]. Inspired by Dai et al. [11], two strategies are adopted to enhance the reliability of LLM-based scoring.
@@
The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies. This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments. We use gemini-2.5-pro-preview-06-05 [3] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
@@
To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses. Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric. The final human score for each metric is calculated as the average of all annotators’ judgments. The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation. The human judgment setting follows the evaluation protocol introduced by Dai et al. [11], which uses pairwise preference comparisons to ensure stable and interpretable annotations. Details are in the appendix.
@@ 5.1 SRPA Architectures
Cascaded SRPAs decompose the overall task into three modular components. First, an ASR model transcribes the user’s spoken input into text. Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role. Finally, a TTS module converts the generated text into speech that matches the target character’s voice. In our implementation, both proprietary and open-source cascaded systems are explored. We utilize a proprietary cascaded system based on the Alibaba Cloud API [1], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2. For the open-source side, we employ whisper-large-v3-turbo [26] for high-quality ASR, LLaMA3.1-8B [15] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [4], CosyVoice2 [14], E2 TTS [16], F5-TTS [10], FireRedTTS [19], GPT-SoVITS [30], Spark-TTS [38], and MaskGCT [40].
@@ 5.2 Evaluation Results
We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity. Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems. Following the same evaluation protocol as Seed-TTS [2], we use the WavLM-large [9] model to extract speaker embeddings and compute cosine similarity scores.
@@ A.1 Text Data Construction
A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [23] and RoleLLM [36]. Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM. The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5). To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics. Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
@@
We use gpt-4.1-2025-04-14 [25] to generate dialogues between users and roles based on role metadata. For some roles with lengthy scripts, the content is divided into segments and processed sequentially. GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions. For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
@@
The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity. First, only dialogues that conform to a clear turn-taking pattern are retained. In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns. To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts. For English data, the all_MiniLM_L6_v2 [28] model is used, and samples with a similarity score below 0.4 are excluded. The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [42] model, with a similarity score threshold of 0.45.
@@ A.2 Role Voice Collection
We first employ a batch processing pipeline based on ffmpeg [18] to extract the raw audio streams from the collected video files. To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format. However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
@@
To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [20] framework. This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation. The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [27] metric. In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3. The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
@@
The extracted audio data contains only the speaker’s index and does not include identity information. To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [12] models to determine the target character’s speaker number based on all transcribed texts. We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
@@
To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity. Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [35] model and calculate the cosine similarity between each sample and all other samples. A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role. Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.

### v4

@@ 1 Introduction
Recently, Role-Playing Agents (RPAs) have emerged as a promising paradigm for achieving personalized interactions and emotional resonance [DBLP:journals/tmlr/Chen00YZSXLYZCL24]. Simulated characters with distinct traits make interactions more engaging in applications such as digital assistants [DBLP:journals/corr/abs-2407-05305], educational tools [DBLP:journals/corr/abs-2409-03512], and interactive storytelling [DBLP:conf/acl/WangPQLZWGGN00024]. However, current research predominantly focuses on the text modality [DBLP:journals/corr/abs-2308-09597, DBLP:conf/emnlp/ShaoLDQ23, DBLP:conf/acl/WangPQLZWGGN00024, DBLP:conf/emnlp/SadeqXKLGM24], overlooking the crucial role of speech in authentic interactions.
Paralinguistic features such as pitch, rhythm, intonation, and timbre are crucial for conveying character personality, emotion, and intent [DBLP:conf/iui/ChangSZJ0K23]. Although preliminary studies have explored role-playing incorporating visual or speech modalities [DBLP:conf/iclr/DaiH0J0025, DBLP:journals/corr/abs-2505-20277], there is currently a lack of systematic frameworks for constructing and evaluating Speech Role-Playing Agents (SRPAs). SRPAs are systems capable of engaging in character-driven spoken dialogue, exhibiting consistent vocal traits, expressive prosody, and coherent persona behavior across interactions. In particular, large-scale, high-quality datasets and evaluation benchmarks for modeling and assessing character consistency, prosodic variation, and interaction capabilities remain scarce.
@@ 2 Related Work
Recent research has increasingly explored the role-playing capabilities of Large Language Models (LLMs), with most efforts focusing on the text modality. ChatHaruhi [DBLP:journals/corr/abs-2308-09597] introduced a retrieval-augmented framework for character-consistent dialogue, while RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024], Character-LLM [DBLP:conf/emnlp/ShaoLDQ23], and RoleFact [DBLP:conf/emnlp/SadeqXKLGM24] proposed benchmarks and profile-grounded generation methods to improve persona simulation and mitigate hallucinations. Beyond text, MMRole [DBLP:conf/iclr/DaiH0J0025] extended role-playing to the visual domain with a multimodal training and evaluation framework, and OmniCharacter [DBLP:journals/corr/abs-2505-20277] demonstrated the value of speech for real-time, voice-aware character interaction.
To assess role-playing capabilities, benchmarks such as RoleEval [DBLP:journals/corr/abs-2312-16132], RoleInteract [DBLP:journals/corr/abs-2403-13679], SocialBench [DBLP:conf/acl/ChenCYXXSQLZH24], and CharacterEval [DBLP:conf/acl/TuFTSSGY24] evaluated behavioral consistency and persona fidelity, while INCHARACTER [DBLP:conf/acl/WangXHYXGTFL0CL24] and ECHO [DBLP:journals/corr/abs-2404-13957] introduced human-aligned, psychometric evaluation paradigms. These advances underscore the importance of role-playing as a core dimension of LLM alignment and interactivity. However, existing efforts primarily target text-based or visual modalities, leaving a significant gap in the systematic evaluation of speech-based role-playing, where vocal expression is crucial for realism, emotional depth, and user engagement.
@@
In recent years, speech-interactive LLMs have emerged as a central topic in human-computer interaction, with mainstream approaches typically falling into two paradigms: cascaded and end-to-end. Cascaded systems perform speech dialogue by sequentially combining automatic speech recognition (ASR), text-based LLMs, and text-to-speech synthesis (TTS). This modularity has led to their wide adoption in systems such as ChatGPT and AudioGPT [DBLP:conf/aaai/HuangLYSCYWHHLR24]. Modern TTS models like F5TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], E2-TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], Spark-TTS [DBLP:journals/corr/abs-2503-01710], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025], and XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24] have significantly improved output quality, expressiveness, and controllability. However, cascaded pipelines often suffer from error propagation and limited cross-modal context integration.
End-to-end systems aim to unify speech perception and generation within a single model. Early efforts like SpeechGPT [DBLP:conf/emnlp/ZhangLZZWZQ23] incorporated discrete speech units into LLM vocabularies, while later models such as AudioPaLM [DBLP:journals/corr/abs-2306-12925] utilized vocoder-based representations for continuous speech modeling. Recent advances, including Moshi [DBLP:journals/corr/abs-2410-00037], LLaMA-Omni [DBLP:conf/iclr/FangGZMZ025], and Qwen2.5-Omni [DBLP:journals/corr/abs-2503-20215], have pushed the boundaries further with dual-modality training, streaming support, and low-latency interactive response, enabling more natural, expressive, and efficient spoken interactions.
@@ 3.1 Text Data Construction
We curate 98 diverse roles, including 18 from ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and 80 from RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024]. The selection balances language, gender, and personality, focusing on characters with distinctive vocal traits to enrich speech modeling.
@@
Using the extracted metadata as a guide, we employ gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate conversations. For each of the 98 roles, we produce approximately 800 single-turn and 800 multi-turn dialogues, creating a substantial initial pool of text data.
@@
The generated data undergoes a multi-stage filtering process. First, we ensure contextual relevance by computing semantic similarity between dialogues and their source metadata, using all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] for English and text2vec-bge-large-chinese [text2vec] for Chinese, discarding samples with scores below 0.4 for English and 0.45 for Chinese. To eliminate redundancy, we remove near-duplicate dialogues using similarity thresholds of 0.9 for English and 0.85 for Chinese.
@@ 3.2 Role Voice Collection
To acquire authentic reference voices, we collect audio from sources like the game Genshin Impact (with permission) and other audiovisual works. All audio is anonymized and fragmented into brief utterances to comply with fair use. Raw streams are extracted using ffmpeg [ffmpeg] and converted to mono 16 kHz WAV format.
@@
We use the open-source Emilia framework [DBLP:conf/slt/HeSWLGHLYLSWCZW24] to process raw audio, including source separation, speaker diarization, and voice activity detection. This yields clean 3–10 second single-speaker clips, transcribed by whisper-medium-v3 [DBLP:conf/icml/RadfordKXBMS23] and scored with DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22].
@@
The diarization process outputs speaker indices without explicit identity information. To attribute the audio to the correct characters, we use LLMs (gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437]) to analyze transcribed text. The models’ integrated results are then manually verified.
@@
To select the most representative audio sample for each role, we rank all of their voice clips based on vocal similarity. We use the CAM++ model [DBLP:conf/interspeech/WangZCC023] to extract a speaker embedding for each clip. The clip with the highest average cosine similarity to all other clips from the same character is selected as the final reference audio, as it best captures the character’s quintessential vocal profile.
@@ 4.3 Evaluation Procedure
There is a lack of widely recognized objective metrics for the evaluation of role speech output, and direct scoring by LLMs often exhibits instability and inconsistency [DBLP:conf/acl/WangLCCZLCKLLS24]. Inspired by DBLP:conf/iclr/DaiH0J0025, two strategies are adopted to enhance the reliability of LLM-based scoring.
@@
The final evaluation score is calculated as the ratio of the test speech score to the reference score, which standardizes the results across different prompts and reduces variance caused by prompt discrepancies. This two-step, reference-based evaluation framework is designed to improve scoring stability and promote better alignment with human judgments. We use gemini-2.5-pro-preview-06-05 [DBLP:journals/corr/abs-2312-11805] to assign scores to each response according to the predetermined evaluation protocol, with scores ranging from 1 to 10.
@@
To verify the reliability of our automatic scoring method, we randomly select 20 dialogue samples and, for each sample, randomly choose two model-generated responses. Four independent human annotators conduct pairwise comparisons of these responses based on multiple evaluation metrics, determining which response performs better for each metric. The final human score for each metric is calculated as the average of all annotators’ judgments. The results show a strong correlation between human evaluation and automatic scoring, confirming that our scoring method aligns closely with human judgments and provides a reliable foundation for large-scale automated evaluation. The human judgment setting follows the evaluation protocol introduced by DBLP:conf/iclr/DaiH0J0025, which uses pairwise preference comparisons to ensure stable and interpretable annotations. Details are in the appendix.
@@ 5.1 SRPA Architectures
Cascaded SRPAs decompose the overall task into three modular components. First, an ASR model transcribes the user’s spoken input into text. Next, a LLM generates a contextually appropriate and character-specific text-based response that aligns with the specified character’s identity, using both the transcription and system prompts defining the role. Finally, a TTS module converts the generated text into speech that matches the target character’s voice. In our implementation, both proprietary and open-source cascaded systems are explored. We utilize a proprietary cascaded system based on the Alibaba Cloud API [alibabacloud], integrating Paraformer Realtime ASR v2, Qwen-Plus-Character, and CosyVoice-v2. For the open-source side, we employ whisper-large-v3-turbo [DBLP:conf/icml/RadfordKXBMS23] for high-quality ASR, LLaMA3.1-8B [DBLP:journals/corr/abs-2407-21783] for flexible and powerful text generation, and multiple state-of-the-art zero-shot timbre cloning TTS models for natural and expressive speech synthesis, including XTTS [DBLP:conf/interspeech/CasanovaDGGGHAM24], CosyVoice2 [DBLP:journals/corr/abs-2412-10117], E2 TTS [DBLP:conf/slt/EskimezWTLTXYZTTLZK24], F5-TTS [DBLP:journals/corr/abs-2410-06885], FireRedTTS [DBLP:journals/corr/abs-2409-03283], GPT-SoVITS [gpt-sovits], Spark-TTS [DBLP:journals/corr/abs-2503-01710], and MaskGCT [DBLP:conf/iclr/WangZLZGZZZZ025].
End-to-end SRPAs generate speech responses directly from spoken input, guided by the user’s spoken query and a textual prompt of the character’s identity. Both the proprietary gpt-4o-audio-preview-2024-12-17 [DBLP:journals/corr/abs-2410-21276] and the open-source Qwen2.5-Omni-7B [DBLP:journals/corr/abs-2503-20215] are evaluated in our study.
@@ 5.2 Evaluation Results
We further analyze the performance of SRPAs with voice cloning capability in terms of speaker similarity. Since end-to-end models do not support speaker cloning, we evaluate speaker similarity only for cascade systems. Following the same evaluation protocol as Seed-TTS [DBLP:journals/corr/abs-2406-02430], we use the WavLM-large [DBLP:journals/jstsp/ChenWCWLCLKYXWZ22] model to extract speaker embeddings and compute cosine similarity scores.
@@ A.1 Text Data Construction
A total of 98 representative roles are meticulously curated for SpeechRole-Data, selected from two high-quality datasets: ChatHaruhi [DBLP:journals/corr/abs-2308-09597] and RoleLLM [DBLP:conf/acl/WangPQLZWGGN00024]. Specifically, 18 roles from games, movies, and television series are chosen from ChatHaruhi, while 80 roles from animation, movies, and television series are drawn from RoleLLM. The selected roles are balanced in terms of language, gender, and personality traits, thereby ensuring diversity in both vocal expression and character identity (see Figure 5). To facilitate rich and character-consistent speech modeling, particular emphasis is placed on roles with distinctive vocal characteristics. Table 5 and Table 6 summarizes the attributes of all roles, including their categories, languages, genders, source materials, and dataset splits.
@@
We use gpt-4.1-2025-04-14 [DBLP:journals/corr/abs-2303-08774] to generate dialogues between users and roles based on role metadata. For some roles with lengthy scripts, the content is divided into segments and processed sequentially. GPT-4.1 is required to provide relevant script segments while generating dialogues to serve as references for SRPA in answering questions. For each role, 800 single-turn dialogues and 800 multi-turn dialogue data are generated.
@@
The data are filtered through a multi-stage process to ensure structural consistency, contextual relevance, and semantic diversity. First, only dialogues that conform to a clear turn-taking pattern are retained. In these dialogues, the user initiates with a question, and the role responds, with each party speaking in alternating turns. To ensure alignment between the dialogue and its corresponding context, semantic embeddings are computed for both texts. For English data, the all_MiniLM_L6_v2 [DBLP:conf/emnlp/ReimersG19] model is used, and samples with a similarity score below 0.4 are excluded. The same procedure is applied to Chinese data using the text2vec-bge-large-chinese [text2vec] model, with a similarity score threshold of 0.45.
@@ A.2 Role Voice Collection
We first employ a batch processing pipeline based on ffmpeg [ffmpeg] to extract the raw audio streams from the collected video files. To ensure consistency, all audio is converted to the standard mono 16 kHz WAV format. However, since these recordings often contain background music, noise, and overlapping speech, additional processing is required.
@@
To obtain clean audio segments from specific speakers, we employ a structured pipeline based on the open-source Emilia [DBLP:conf/slt/HeSWLGHLYLSWCZW24] framework. This process includes source separation to remove background noise, speaker separation to distinguish individual speakers, and voice activity detection for fine-grained segmentation. The resulting audio segments, each lasting between 3 and 10 seconds and containing only a single speaker, are well-suited for subsequent modeling.
Finally, we use an automatic speech recognition model to transcribe each segmented audio fragment and evaluate speech quality using the DNSMOS P.835 OVRL [DBLP:conf/icassp/ReddyGC22] metric. In order to ensure the quality of speech data, we filter out speech with DNMOS less than 3. The resulting structured dataset provides, for each audio entry, the following information: start and end times, transcription, speaker index, language, and DNSMOS-based audio quality score.
@@
The extracted audio data contains only the speaker’s index and does not include identity information. To identify the speaker, we use both gpt-4.1-2025-04-14 and DeepSeek-V3-0324 [DBLP:journals/corr/abs-2412-19437] models to determine the target character’s speaker number based on all transcribed texts. We integrate the results from both models and verify them manually to obtain the speech corresponding to the target character.
@@
To select the most representative audio sample as a reference for subsequent role-playing, we rank all speech samples of each speaker based on similarity. Specifically, we extract the speaker feature vector of each audio sample using the CAM++ [DBLP:conf/interspeech/WangZCC023] model and calculate the cosine similarity between each sample and all other samples. A higher average similarity score indicates that the audio sample better represents the overall speaker characteristics of the role. Finally, we sort all audio samples for each role in descending order according to their similarity scores and select the top-ranked sample as the reference audio.